OpenStudy (anonymous):

Richard spent 21.2% of his savings account during the summer and now has $6,030.72 remaining in the account. What was the amount in his savings account at the beginning of the summer? Please show all work

OpenStudy (anonymous):

100 - 21.2 = 6030.72

OpenStudy (anonymous):

He started with x money He spent 21.2 percent and now has 6030.72 left This could be written as x*(1-0.212)=6030.72 Divide both sides by 0.788 x=7653.2 To check answer, 7653.2-(7653.2*.212)=6030.72 Answer:Richard's savings at the beginning of the summer were $7653.29

OpenStudy (anonymous):

?

OpenStudy (anonymous):

x78.8/100=6030.72 x = 6030.72*100/78.8=

OpenStudy (anonymous):

7653.198